OpenStudy (anonymous):

Find x if y=1/2. 4x/y=2(x+1) A.3 B.–3 C. 1/3 D.0

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

4(x) / (1/2) = 2(x + 1) = 8x = 2x + 2 6x = 2 x = 1/3

OpenStudy (anonymous):

Thank you!

OpenStudy (blurbendy):

np